P 147 of 1065 9. The composition of the Fingroup Fund portfolio is as follows: Price Shares 200,000 300,000 400,000 600,000 Stock S35 40 20 25 manager currently totals $30,000. There are 4 million shares outstanding. What is the net asset value of the fund? Reconsider the Fingroup Fund in the previous problem. If during the year the portfolio manager sells all of the holdings of stock D and replaces it with 200,000 shares of stock E at $50 per share and 200,000 shares of stock F at $25 per share, what is the portfolio turnover rate? 10.
Answer #1
09) NAV = (200000*35+300000*40+400000*20+600000*25-30000)/4000000 = $       10.49
10) Portfolio turnover rate = The total purchases or total sales, whichever is lower/average of the net assets.
Total purchases = 200000*50+200000*25 = 15000000
Total sales = 200000*25 = 15000000
Net assets at year beginning = (200000*35+300000*40+400000*20+600000*25) = 42000000
Net assets at year end = (200000*35+300000*40+400000*20+600000*25-30000) = 41970000
Average net assets = (41970000+42000000)/2 = 41985000
Portfolio turnover rate = 15000000/41985000 = 35.73%